Q(2)A random variable is defined as Construct the probability distribution function

of X and find the standard deviation of X.


1
Expert's answer
2021-08-10T11:51:45-0400

The probability distribution for discrete random variable X can be represented by the formula

p(x)=P(X=x) for all x.

"V(x)=E[(X-\mu)^2]"

Standard deviation of "X=\sqrt{V(x)}"

"=\sqrt{E[(X-\mu)^2]}"